# Unified Multimodal Understanding and Generation Models: Advances, Challenges, and Opportunities

### v4

@@ 4.5 Other Text+Image-to-Image Datasets
General-Bench [314] $>$700 tasks Ultra Large-scale Eval Varied by Task Arxiv2025
@@
ComplexBench-Edit [340] $<$1K List of Edits Chain-dependent Editing Eval Image + Instruction $\rightarrow$ Image Arxiv2025
@@ 5.2 Evaluation on Image Generation
To support large-scale concept diversity, EvalMuse-40K [355] provides 40K crowdsourced prompts focusing on nuanced concept representation, and HEIM [321] identifies 12 aspects, including text-image alignment, image quality, aesthetics, originality, reasoning, knowledge, bias, toxicity, fairness, robustness, multilinguality and efficiency. Considering practical needs, FlashEval [356] shrinks the large-scale evaluation set into diverse smaller ones through iterative search to accelerate the benchmark testing. MEMO-Bench [357] introduces a comprehensive benchmark for evaluating the emotional understanding and expression capabilities of T2I models and MLLMs. ConceptMix [325] evaluates text-to-image models’ compositional generation ability by sampling k-tuples of visual concepts to construct prompts and automatically verifying concept presence in the resulting images using a strong visual language model. TIFA [318] offers a fine-grained benchmark for evaluating text-to-image faithfulness via visual question answering generated from prompts. To enrich dependencies in question generation for VQA-based evaluation of image–prompt alignment, DSG-1k [323] refines its questions using a multi-level semantic graph. MMIG-Bench [328] introduced a multi-dimensional assessment framework that rigorously examines text-to-image generation models.
@@ 6 Challenges and Opportunities on Unified Models
Currently, at its rudimentary stage, unified multimodal models face several significant challenges that should be addressed to achieve robust and scalable understanding and generation capabilities. First, the high dimensionality of visual and textual data leads to extremely long token sequences. Efficient tokenization and compression strategies are essential to reduce memory and computation costs while preserving representational fidelity. Second, cross-modal attention becomes a performance bottleneck as image resolution and context length increase. Scalable alternatives such as sparse or hierarchical attention mechanisms may potentially mitigate this issue. Third, pretraining datasets often include noisy or biased image–text pairs, particularly for complex image compositions and interleaved image-text data. Reliable data filtering, debiasing, and synthesizing are crucial to ensure fairness and robustness. Fourth, evaluation protocols are typically designed for single tasks in isolation. There is a growing need for comprehensive benchmarks that assess both understanding and generation in an integrated manner, especially for sophisticated tasks such as image editing and interleaved image-text generation. Apart from the issues and challenges on architecture, data, and evaluation, applying chain-of-thought (CoT) reasoning and reinforcement learning (RL) techniques into unified MLLM models to improve both interpretability and performance [359] is also worth exploring. CoT can guide the model to generate intermediate reasoning steps, which are particularly beneficial for complex visual-question answering or image-conditioned generation. Meanwhile, RL can be used to optimize long-horizon objectives such as factual consistency, user satisfaction, or task success rate beyond token-level likelihoods. Moreover, exploring the demographic and social biases of existing unified MLLM models [360] is an important topic to ensure responsible deployment. As these models become increasingly capable across diverse modalities and tasks, unintentional amplification of cultural stereotypes, gender bias, or geographic imbalances embedded in pretraining data may result in harmful outputs. Future work should investigate effective fairness-aware training pipelines. Finally, enabling personalized knowledge-driven generation within unified MLLMs [361] is an emerging and important direction. Personalized models aim to incorporate user-provided concepts—such as specific objects, characters, or styles—into the model’s understanding and generation capabilities. However, current approaches often treat understanding and generation separately, using distinct concept embeddings for each task. This separation limits the model’s ability to generalize to compositional prompts that require implicit knowledge, such as generating ”$<$bo$>$ wearing its hat” without explicitly describing the hat. Unifying personalized understanding and generation under a shared modeling framework would allow better semantic grounding and contextual generalization.

### v5

@@ 4.3 Image Editing Datasets
  • GPT-Image-Edit-1.5M [280]: GPT-IMAGE-EDIT-1.5M is a large-scale image editing dataset containing over 1.5 million high-quality instruction-guided image editing triplets. It is constructed by leveraging the powerful capabilities of GPT-4o to systematically unify and refine three existing datasets: OmniEdit, HQ-Edit, and UltraEdit. The core methodology involves regenerating output images to enhance visual quality and instruction alignment, as well as selectively rewriting prompts to improve semantic clarity. This process results in a high-fidelity corpus designed to bridge the gap between proprietary and open-source instruction-guided image editing models.
  • X2Edit [281]: X2Edit is a large-scale and comprehensive image editing dataset with 3.7 million samples, designed to be balanced across 14 diverse editing tasks. It is constructed through an automated pipeline that first uses a VLM to generate task-aware instructions, which are then executed by industry-leading and expert generative models to produce the edited images. To overcome the quality and balance issues in existing open-source resources, all generated pairs undergo a final, rigorous filtering stage based on multiple scoring mechanisms to ensure high fidelity and accuracy.
@@ 4.5 Other Text+Image-to-Image Datasets
  • Echo-4o-Image (Multi-Reference) [267]: This component of the dataset addresses the scarcity of structured, multi-input generation tasks in natural image collections. It provides 73K synthetic samples for “Multi-to-one” generation, which are explicitly designed with diverse instructions and rich reference information. This controlled synthesis offers a more targeted and varied training source for multi-reference image composition compared to alternatives like sampling from video frames.
@@
General-Bench [324] >>700 tasks Ultra Large-scale Eval Varied by Task Arxiv2025
@@
OneIG-Bench [339] $\sim$2k Prompts Comprehensive Eval Text Prompt $\rightarrow$ Image Arxiv2025
WISE [340] 1k Prompts World Knowledge Eval Text Prompt $\rightarrow$ Image Arxiv2025
CVTG-2K [341] 2k Prompts Multi-region Visual Text Eval Text Prompt $\rightarrow$ Image Arxiv2025
WorldGenBench [342] 1072 Prompts World Knowledge Eval Text Prompt $\rightarrow$ Image Arxiv2025
@@
ComplexBench-Edit [354] <<1K List of Edits Chain-dependent Editing Eval Image + Instruction $\rightarrow$ Image Arxiv2025
@@ 5.2 Evaluation on Image Generation
To support large-scale concept diversity, EvalMuse-40K [369] provides 40K crowdsourced prompts focusing on nuanced concept representation, and HEIM [331] identifies 12 aspects, including text-image alignment, image quality, aesthetics, originality, reasoning, knowledge, bias, toxicity, fairness, robustness, multilinguality and efficiency. Considering practical needs, FlashEval [370] shrinks the large-scale evaluation set into diverse smaller ones through iterative search to accelerate the benchmark testing. MEMO-Bench [371] introduces a comprehensive benchmark for evaluating the emotional understanding and expression capabilities of T2I models and MLLMs. ConceptMix [335] evaluates text-to-image models’ compositional generation ability by sampling k-tuples of visual concepts to construct prompts and automatically verifying concept presence in the resulting images using a strong visual language model. TIFA [328] offers a fine-grained benchmark for evaluating text-to-image faithfulness via visual question answering generated from prompts.
To enrich dependencies in question generation for VQA-based evaluation of image–prompt alignment, DSG-1k [333] refines its questions using a multi-level semantic graph. MMIG-Bench [338] introduced a multi-dimensional assessment framework that rigorously examines text-to-image generation models. OneIG-Bench [339] introduces a comprehensive fine-grained evaluation framework for text-to-image models across more dimensions.  [340, 342] evaluate text-to-image models’ world knowledge understanding, which emphasizes semantic consistency, realism, and aesthetics. CVTG-2K [341] evaluates visual-text generation on complex multi-region layouts, diverse text attributes, and fine-grained positioning.
@@ 6 Challenges and Opportunities on Unified Models
Currently, at its rudimentary stage, unified multimodal models face several significant challenges that should be addressed to achieve robust and scalable understanding and generation capabilities. First, the high dimensionality of visual and textual data leads to extremely long token sequences. Efficient tokenization and compression strategies are essential to reduce memory and computation costs while preserving representational fidelity. Second, cross-modal attention becomes a performance bottleneck as image resolution and context length increase. Scalable alternatives such as sparse or hierarchical attention mechanisms may potentially mitigate this issue. Third, pretraining datasets often include noisy or biased image–text pairs, particularly for complex image compositions and interleaved image-text data. Reliable data filtering, debiasing, and synthesizing are crucial to ensure fairness and robustness. Fourth, evaluation protocols are typically designed for single tasks in isolation. There is a growing need for comprehensive benchmarks that assess both understanding and generation in an integrated manner, especially for sophisticated tasks such as image editing and interleaved image-text generation. Apart from the issues and challenges on architecture, data, and evaluation, applying chain-of-thought (CoT) reasoning and reinforcement learning (RL) techniques into unified MLLM models to improve both interpretability and performance [373] is also worth exploring. CoT can guide the model to generate intermediate reasoning steps, which are particularly beneficial for complex visual-question answering or image-conditioned generation. Meanwhile, RL can be used to optimize long-horizon objectives such as factual consistency, user satisfaction, or task success rate beyond token-level likelihoods. Moreover, exploring the demographic and social biases of existing unified MLLM models [374] is an important topic to ensure responsible deployment. As these models become increasingly capable across diverse modalities and tasks, unintentional amplification of cultural stereotypes, gender bias, or geographic imbalances embedded in pretraining data may result in harmful outputs. Future work should investigate effective fairness-aware training pipelines. Finally, enabling personalized knowledge-driven generation within unified MLLMs [375] is an emerging and important direction. Personalized models aim to incorporate user-provided concepts—such as specific objects, characters, or styles—into the model’s understanding and generation capabilities. However, current approaches often treat understanding and generation separately, using distinct concept embeddings for each task. This separation limits the model’s ability to generalize to compositional prompts that require implicit knowledge, such as generating ”<<bo>> wearing its hat” without explicitly describing the hat. Unifying personalized understanding and generation under a shared modeling framework would allow better semantic grounding and contextual generalization.